# Spectral Convolutional Conditional Neural Processes

### v5

@@ Appendix A Discretization Sensitivity of the DFT
To illustrate the sensitivity of the discrete Fourier transform (DFT) to grid shape—and how changes in discretization can disturb predictive performance—we consider a simple one-dimensional example. Let $0<\Delta\ll 1$ denote a fixed discretization resolution, interpreted as the spacing between successive points. Define the grid $\mathcal{G}_{1}=\left(x_{0},\dots,x_{m_{1}-1}\right)$ as a uniform discretization of the interval $[0,1]$ with $m_{1}=\lfloor\frac{1}{\Delta}\rfloor$ and ordered such that $x_{k-1}<x_{k}$ for all $k\in\{1,\dots,m_{1}-1\}$. Let the sampled values of $\varphi_{e}[\mathcal{D}_{c}]$ at $\mathcal{G}_{1}$ be denoted by $\left(\varphi_{e}\left[\mathcal{D}_{c}\right](x)\right)_{x\in\mathcal{G}_{1}}=\left(\varphi_{e}[\mathcal{D}_{c}](x_{0}),\;\dots\;,\varphi_{e}[\mathcal{D}_{c}](x_{m_{1}-1})\right)$. The DFT of $\varphi_{e}[\mathcal{D}_{c}]$ over this grid is given by:
@@
Thus, the DFT yields Fourier coefficients at the normalized frequencies $\hat{\Xi}_{1}=\left(\frac{k}{m_{1}}\right)_{k=0}^{m_{1}-1}$. As described in Section 2.1, the FNO implicitly ties its kernel spectrum parameterization to the set $\hat{\Xi}_{1}$ encountered during training. Now suppose we wish to evaluate the model on a larger domain, say $[0,2]$, using the same resolution $\Delta$. Then the corresponding grid $\mathcal{G}_{2}$ has size $m_{2}=\lfloor\frac{2}{\Delta}\rfloor$, and the DFT computes coefficients at normalized frequencies $\hat{\Xi}_{2}=\left(\frac{k}{m_{2}}\right)_{k=0}^{m_{2}-1}$. Since $m_{2}\geq 2m_{1}$, we have $|\hat{\Xi}_{2}|\geq 2|\hat{\Xi}_{1}|$. Let $1\leq k_{\text{max}}\leq m_{1}$ denote the number of frequency modes retained during training. Clearly, the first $k_{\text{max}}$ frequencies of $\hat{\Xi}_{2}$ do not align with those of $\hat{\Xi}_{1}$, resulting in a mismatch.
@@ Appendix B Experimental Details
All implementations are written in PyTorch [paszke2019pytorch] and publicly available at https://github.com/peiman-m/SConvCNP. We used a single NVIDIA A100 GPU with 40 GB of memory for all the computations. Our code is based on the implementations of ashman2024translation and bruinsma2023autoregressive.
@@ B.1.1 Model Architectures
This section details the architectures of all CNPs variants used in our experiments. Each model outputs a Gaussian predictive distribution parameterized by a mean and a pre-softplus [dugas2000incorporating] scale. The pre-softplus value is passed through a softplus transformation, and a minimum noise term of $10^{-6}$ is then added to the resulting scale. The Gaussian factorizes across query points—as is standard for CNPs (see Section 2.2)—and, for multidimensional outputs, across output dimensions as well. Thus, predictions assume independence both across query locations and across components of each output vector. Unless otherwise noted, all nonlinearities use ReLU activations [nair2010rectified].
@@
The CNP encodes each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ using separate input and output pathways. The input $x_{c,k}$ and output $y_{c,k}$ are first passed through distinct MLPs with two hidden layers each of dimension 256, producing 256-dimensional representations $\varepsilon^{(x)}_{c,k}$ and $\varepsilon^{(y)}_{c,k}$. Their concatenation $\varepsilon_{c,k}=[\varepsilon^{(x)}_{c,k},\varepsilon^{(y)}_{c,k}$] is then processed by a deeper MLP with six hidden layers of width 256 and a final 256-dimensional output layer, yielding an embedding for each context pair. Context embeddings $\varepsilon_{c,k}$ are averaged to form a single aggregated representation $\varepsilon_{c}$ of dimension 256. For prediction, each query input $x_{q,k}$ is concatenated with the context representation $\varepsilon_{c}$ and passed through a decoder MLP consisting of six hidden layers of dimension 256. The decoder’s final output layer has dimensionality $2d_{y}$, parameterizing the mean and the log-scale parameters of the Gaussian predictive distribution.
@@
The AttCNP implementation follows the deterministic architecture introduced by kim2019attentive. The initial encoding stage mirrors that of the CNP: each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ is processed by two separate MLPs (each with two hidden layers of width 256), producing embeddings $\varepsilon^{(x)}_{c,k}$ and $\varepsilon^{(y)}_{c,k}$. These are concatenated to form $\varepsilon_{c,k}=[\varepsilon^{(x)}_{c,k},\varepsilon^{(y)}_{c,k}$], which is subsequently passed through an additional two-layer MLP with hidden width 256. Departing from the CNP, the AttCNP applies self-attention to the set of context embeddings. We use two layers of multi-head self-attention, each with 8 heads (head dimension 32), a feedforward subnetwork with hidden width 256, residual connections, and pre-layer normalization for both the attention and feedforward blocks. The input and output dimensionality of each attention layer is fixed at 256. Query representations $\varepsilon_{q,k}$ are computed using a single layer of multi-head cross-attention between the (attended) context embeddings and the query embeddings, with the same configuration as above (8 heads of dimension 32, feedforward width 256, residual connections, and pre-normalization). The decoder is an MLP with six hidden layers of width 256. Its final layer outputs $2d_{y}$ units, corresponding to the mean and log-scale parameters of a Gaussian predictive distribution.
@@
where the final singleton “1” serves as a density flag indicating that the observation $y_{c,k}$ at $x_{c,k}$ is available. For each query location $x_{q,k}$, we instead construct
$x_{q,k},\,\mathbf{0},\,0],$
where the dummy zero-vector matches the shape of $y_{c,k}$ and the final “0” indicates the absence of an observation [nguyen2022transformer, ashman2024translation]. All tokens are then passed through a shared two-layer MLP (width 256), producing initial embeddings $\varepsilon_{c,k}$ and $\varepsilon_{q,k}$.
@@
Because queries never attend to one another, feng2022latent introduced the Efficient Query TNP, which we adopt here. This variant first applies self-attention to context tokens to produce updated context embeddings, then processes each query using cross-attention over these updated contexts. The resulting two-branch structure (contexts processed twice, queries once) reduces the overall complexity to
@@
while for each query location $x_{q,k}$ we use the token
@@
All tokens are passed through a shared two-layer MLP with width 256 to produce initial embeddings $\varepsilon_{c,k}$ and $\varepsilon_{q,k}$.
The TETNP replaces the standard multi-head attention mechanism with the translation-equivariant attention proposed by ashman2024translation. For each attention head, we first compute the matrix of pairwise scaled dot products between token embeddings, analogous to conventional attention but independent of absolute input locations. In parallel, we compute the matrix of pairwise differences between token locations. We then concatenate the scaled dot-product scores from all heads with the corresponding pairwise location differences. This augmented pairwise similarity representation is processed by an MLP—acting as an implicit kernel—with two hidden layers of width 256 and an output dimension equal to the number of attention heads, yielding the final translation-equivariant attention logits.
@@
The functional embedding in equation 3 is evaluated on this grid. The Gaussian kernels used in this embedding are initialized with per-dimension length scales set to twice the grid resolution, i.e., $2/64$. For multi-dimensional inputs and outputs, the model uses separate length scales for each dimension. Following bruinsma2023autoregressive, the embedding is further normalized using a density channel, defined as
@@
Each grid point is processed independently by an MLP with two hidden layers of width 128. The resulting features are passed to a CNN based on a U-Net architecture [ronneberger2015u], consisting of six residual convolutional blocks (kernel size 11, stride 2, 128 channels) in the encoder, followed by a symmetric sequence of six transposed convolutional blocks in the decoder, with skip connections following the design of bruinsma2023autoregressive. Since the U-Net downsamples the spatial resolution by a factor of 64, we ensure that the constructed grid size is divisible by 64. For the one-dimensional benchmarks considered in this work, this is achieved by symmetrically enlarging the discretization interval when necessary so that the CNN output aligns exactly with the input grid.
@@
The SConvCNP replaces the ConvCNP’s U-Net backbone with a U-shaped Fourier Neural Operator (FNO) architecture [li2020fourier, rahman2022u]. A residual Fourier block is denoted
@@
The representation from equation 5, augmented with positional encodings (grid coordinates), is first processed by an MLP with one hidden layer of width 64 and GELU activations [hendrycks2016gaussian]. Its output is then passed through the following sequence of residual Fourier blocks:
@@ B.1.2 Data and Experimental Setup
We evaluate our methods on four families of synthetic 1D regression tasks, each defined by a distinct stochastic generative process: a Gaussian process (GP) with a Matérn–5/2 kernel, a GP with a periodic kernel, a sawtooth-wave generator, and a square-wave generator. GP tasks are sampled using GPyTorch [gardner2018gpytorch]. All processes include independent Gaussian observation noise.
@@
Each model is trained for 500 epochs using AdamW [loshchilov2017decoupled] with a learning rate of $5\times 10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$ [pascanu2013difficulty]. Each epoch consists of 1000 iterations, and every iteration processes a batch of 16 tasks, yielding a total of 8 million on-the-fly sampled training tasks. For each batch, the numbers of context and query points are independently drawn as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$ and $|\mathcal{I}_{q}|\sim\mathcal{U}[5,25)$. These values are shared across all tasks in the batch. Input locations are sampled uniformly and independently from [3,3)[-3,3) for each task.
@@ B.2.2 Data and Experimental Setup
We generate data from a stochastic variant of the Lotka–Volterra predator–prey system [Lotka1910contribution, Volterra1926variazioni], following the formulation of bruinsma2023autoregressive. Let $U_{t}$ and $V_{t}$ denote the prey and predator populations at time $t$. Their dynamics evolve according to
@@
Models are trained for 500 epochs using AdamW with a learning rate of $10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$. Each epoch consists of 1,000 iterations, each processing a batch of 16 tasks, yielding a total of 8 million on-the-fly sampled training tasks. For every batch, the numbers of context and query points are drawn independently as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$ and $|\mathcal{I}_{q}|\sim\mathcal{U}[5,25)$; the sampled values are shared across all tasks within the batch.
@@ B.3.2 Data and Experimental Setup
We use the California traffic-flow dataset from LargeST [liu2023largest], a large-scale benchmark for traffic forecasting, available at https://www.kaggle.com/datasets/liuxu77/largest. The dataset contains traffic-flow measurements from 8,600 loop-detector sensors deployed across California’s highway network, collected at 5-minute intervals between 2017 and 2021. In our experiments, we restrict attention to data from the year 2020.
@@
Models are trained for 100 epochs using AdamW with a learning rate of $10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$. Each epoch consists of approximately 4,160 iterations, each processing a batch of 32 tasks. For every batch, the numbers of context and query points are drawn independently as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,25)$ and $|\mathcal{I}_{q}|\sim\mathcal{U}[5,25)$; the sampled values are shared across all tasks within the batch. A task is formed by independently and uniformly sampling $|\mathcal{I}_{c}|+|\mathcal{I}_{q}|$ time steps from a window without replacement. Training tasks are sampled on-the-fly from training windows. For validation and testing, we pre-generate fixed sets of tasks:
  • Validation: 22,178 tasks (batched in 32), with $|\mathcal{I}_{q}|=256$ and $|\mathcal{I}_{c}|$ sampled as in training.
  • Test: 66,586 tasks (batched in 32), with $|\mathcal{I}_{q}|=50$ and $|\mathcal{I}_{c}|$ sampled as in training.

### v6

@@ Appendix A Discretization Sensitivity of the DFT
To illustrate the sensitivity of the discrete Fourier transform (DFT) to the underlying discretization—and how changes in grid geometry can degrade predictive performance—we consider a simple one-dimensional example. Let $0<\Delta\ll 1$ denote a fixed discretization resolution, interpreted as the spacing between adjacent grid points. Define the grid
$\mathcal{G}_{1}=(x_{0},\dots,x_{m_{1}-1})$
as a uniform discretization of the interval $[0,1]$, where $m_{1}=\lfloor 1/\Delta\rfloor$ and the points are ordered such that $x_{n}<x_{n+1}$ for all $n\in\{0,\dots,m_{1}-2\}$. Let $h:=\varphi_{e}[\mathcal{D}_{c}]$ denote the encoded latent function, and let its sampled values on $\mathcal{G}_{1}$ be
$\bigl(h(x)\bigr)_{x\in\mathcal{G}_{1}}=\bigl(h(x_{0}),\dots,h(x_{m_{1}-1})\bigr).$
The DFT of $h$ on this grid is given by
@@
This yields Fourier coefficients at the normalized frequencies
$\Xi_{1}=\left(\tfrac{k}{m_{1}}\right)_{k=0}^{m_{1}-1}.$
As discussed in Section 2.1, the Fourier Neural Operator (FNO) implicitly ties its kernel parameterization to the specific frequency set $\Xi_{1}$ encountered during training. Now suppose the trained model is evaluated on a larger spatial domain, e.g. $[0,2]$, while maintaining the same resolution $\Delta$. The corresponding grid
$\mathcal{G}_{2}=(x^{\prime}_{0},\dots,x^{\prime}_{m_{2}-1})$
has size $m_{2}=\lfloor 2/\Delta\rfloor$, and the associated DFT produces coefficients at normalized frequencies
$\Xi_{2}=\left(\tfrac{k}{m_{2}}\right)_{k=0}^{m_{2}-1}.$
Since $m_{2}\geq 2m_{1}$, we have $|\Xi_{2}|\geq 2|\Xi_{1}|$, and, crucially, the frequency locations themselves differ. Let $1\leq k_{\max}\leq m_{1}$ denote the number of Fourier modes retained during training. Even when restricting attention to the lowest $k_{\max}$ modes, the frequencies $\{k/m_{2}\}_{k=0}^{k_{\max}-1}$ do not coincide with $\{k/m_{1}\}_{k=0}^{k_{\max}-1}$. This misalignment induces a spectral mismatch between training and evaluation, illustrating how changes in discretization—even at fixed resolution—can disrupt the learned Fourier parameterization and lead to degraded generalization.
@@ Appendix B Experimental Details
All implementations are written in PyTorch [Paszke et al., 2019] and publicly available at https://github.com/peiman-m/SConvCNP. We used a single NVIDIA A100 GPU with 40 GB of memory for all the computations. Our code is based on the implementations of Ashman et al. [2024a] and Bruinsma et al. [2023].
@@ B.1.1 Model Architectures
This section details the architectures of all CNPs variants used in our experiments. Each model outputs a Gaussian predictive distribution parameterized by a mean and a pre-softplus [Dugas et al., 2000] scale. The pre-softplus value is passed through a softplus transformation, and a minimum noise term of $10^{-6}$ is then added to the resulting scale. The Gaussian factorizes across query points—as is standard for CNPs (see Section 2.2)—and, for multidimensional outputs, across output dimensions as well. Thus, predictions assume independence both across query locations and across components of each output vector. Unless otherwise noted, all nonlinearities use ReLU activations [Nair and Hinton, 2010].
@@
The CNP encodes each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ using separate input and output pathways. The input $x_{c,k}$ and output $y_{c,k}$ are first passed through distinct MLPs with two hidden layers each of dimension 256, producing 256-dimensional representations $\varepsilon^{(x)}_{c,k}$ and $\varepsilon^{(y)}_{c,k}$. Their concatenation $\varepsilon_{c,k}=[\varepsilon^{(x)}_{c,k},\varepsilon^{(y)}_{c,k}$] is then processed by a deeper MLP with six hidden layers of width 256 and a final 256-dimensional output layer, yielding an embedding for each context pair. Context embeddings $\varepsilon_{c,k}$ are averaged to form a single aggregated representation $\varepsilon_{c}$ of dimension 256. For prediction, each query input $x_{q,l}$ is concatenated with the context representation $\varepsilon_{c}$ and passed through a decoder MLP consisting of six hidden layers of dimension 256. The decoder’s final output layer has dimensionality $2d_{y}$, parameterizing the mean and the log-scale parameters of the Gaussian predictive distribution.
@@
The AttCNP implementation follows the deterministic architecture introduced by Kim et al. [2019]. The initial encoding stage mirrors that of the CNP: each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ is processed by two separate MLPs (each with two hidden layers of width 256), producing embeddings $\varepsilon^{(x)}_{c,k}$ and $\varepsilon^{(y)}_{c,k}$. These are concatenated to form $\varepsilon_{c,k}=[\varepsilon^{(x)}_{c,k},\varepsilon^{(y)}_{c,k}$], which is subsequently passed through an additional two-layer MLP with hidden width 256. Departing from the CNP, the AttCNP applies self-attention to the set of context embeddings. We use two layers of multi-head self-attention, each with 8 heads (head dimension 32), a feedforward subnetwork with hidden width 256, residual connections, and pre-layer normalization for both the attention and feedforward blocks. The input and output dimensionality of each attention layer is fixed at 256. Query representations $\varepsilon_{q,l}$ are computed using a single layer of multi-head cross-attention between the (attended) context embeddings and the query embeddings, with the same configuration as above (8 heads of dimension 32, feedforward width 256, residual connections, and pre-normalization). The decoder is an MLP with six hidden layers of width 256. Its final layer outputs $2d_{y}$ units, corresponding to the mean and log-scale parameters of a Gaussian predictive distribution.
@@
where the final singleton “1” serves as a density flag indicating that the observation $y_{c,k}$ at $x_{c,k}$ is available. For each query location $x_{q,l}$, we instead construct
$x_{q,l},\,\mathbf{0},\,0],$
where the dummy zero-vector matches the shape of $y_{c,k}$ and the final “0” indicates the absence of an observation [Nguyen and Grover, 2022, Ashman et al., 2024a]. All tokens are then passed through a shared two-layer MLP (width 256), producing initial embeddings $\varepsilon_{c,k}$ and $\varepsilon_{q,l}$.
@@
Because queries never attend to one another, Feng et al. [2022] introduced the Efficient Query TNP, which we adopt here. This variant first applies self-attention to context tokens to produce updated context embeddings, then processes each query using cross-attention over these updated contexts. The resulting two-branch structure (contexts processed twice, queries once) reduces the overall complexity to
@@
while for each query location $x_{q,l}$ we use the token
@@
All tokens are passed through a shared two-layer MLP with width 256 to produce initial embeddings $\varepsilon_{c,k}$ and $\varepsilon_{q,l}$.
The TETNP replaces the standard multi-head attention mechanism with the translation-equivariant attention proposed by Ashman et al. [2024a]. For each attention head, we first compute the matrix of pairwise scaled dot products between token embeddings, analogous to conventional attention but independent of absolute input locations. In parallel, we compute the matrix of pairwise differences between token locations. We then concatenate the scaled dot-product scores from all heads with the corresponding pairwise location differences. This augmented pairwise similarity representation is processed by an MLP—acting as an implicit kernel—with two hidden layers of width 256 and an output dimension equal to the number of attention heads, yielding the final translation-equivariant attention logits.
@@
The functional embedding in equation 3 is evaluated on this grid. The Gaussian kernels used in this embedding are initialized with per-dimension length scales set to twice the grid resolution, i.e., $2/64$. For multi-dimensional inputs and outputs, the model uses separate length scales for each dimension. Following Bruinsma et al. [2023], the embedding is further normalized using a density channel, defined as
@@
Each grid point is processed independently by an MLP with two hidden layers of width 128. The resulting features are passed to a CNN based on a U-Net architecture [Ronneberger et al., 2015], consisting of six residual convolutional blocks (kernel size 11, stride 2, 128 channels) in the encoder, followed by a symmetric sequence of six transposed convolutional blocks in the decoder, with skip connections following the design of Bruinsma et al. [2023]. Since the U-Net downsamples the spatial resolution by a factor of 64, we ensure that the constructed grid size is divisible by 64. For the one-dimensional benchmarks considered in this work, this is achieved by symmetrically enlarging the discretization interval when necessary so that the CNN output aligns exactly with the input grid.
@@
The SConvCNP replaces the ConvCNP’s U-Net backbone with a U-shaped Fourier Neural Operator (FNO) architecture [Li et al., 2020a, Rahman et al., 2022a]. A residual Fourier block is denoted
@@
The representation from equation 5, augmented with positional encodings (grid coordinates), is first processed by an MLP with one hidden layer of width 64 and GELU activations [Hendrycks and Gimpel, 2016]. Its output is then passed through the following sequence of residual Fourier blocks:
@@ B.1.2 Data and Experimental Setup
We evaluate our methods on four families of synthetic 1D regression tasks, each defined by a distinct stochastic generative process: a Gaussian process (GP) with a Matérn–5/2 kernel, a GP with a periodic kernel, a sawtooth-wave generator, and a square-wave generator. GP tasks are sampled using GPyTorch [Gardner et al., 2018]. All processes include independent Gaussian observation noise.
@@
Each model is trained for 500 epochs using AdamW [Loshchilov and Hutter, 2017] with a learning rate of $5\times 10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$ [Pascanu et al., 2013]. Each epoch consists of 1000 iterations, and every iteration processes a batch of 16 tasks, yielding a total of 8 million on-the-fly sampled training tasks. For each batch, the numbers of context and query points are independently drawn as $n_{c}\sim\mathcal{U}[5,25)$ and $n_{q}\sim\mathcal{U}[5,25)$. These values are shared across all tasks in the batch. Input locations are sampled uniformly and independently from [3,3)[-3,3) for each task.
@@ B.2.2 Data and Experimental Setup
We generate data from a stochastic variant of the Lotka–Volterra predator–prey system [Lotka, 1910, Volterra, 1926], following the formulation of Bruinsma et al. [2023]. Let $U_{t}$ and $V_{t}$ denote the prey and predator populations at time $t$. Their dynamics evolve according to
@@
Models are trained for 500 epochs using AdamW with a learning rate of $10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$. Each epoch consists of 1,000 iterations, each processing a batch of 16 tasks, yielding a total of 8 million on-the-fly sampled training tasks. For every batch, the numbers of context and query points are drawn independently as $n_{c}\sim\mathcal{U}[5,25)$ and $n_{q}\sim\mathcal{U}[5,25)$; the sampled values are shared across all tasks within the batch.
@@ B.3.2 Data and Experimental Setup
We use the California traffic-flow dataset from LargeST [Liu et al., 2023b], a large-scale benchmark for traffic forecasting, available at https://www.kaggle.com/datasets/liuxu77/largest. The dataset contains traffic-flow measurements from 8,600 loop-detector sensors deployed across California’s highway network, collected at 5-minute intervals between 2017 and 2021. In our experiments, we restrict attention to data from the year 2020.
@@
Models are trained for 100 epochs using AdamW with a learning rate of $10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$. Each epoch consists of approximately 4,160 iterations, each processing a batch of 32 tasks. For every batch, the numbers of context and query points are drawn independently as $n_{c}\sim\mathcal{U}[5,25)$ and $n_{q}\sim\mathcal{U}[5,25)$; the sampled values are shared across all tasks within the batch. A task is formed by independently and uniformly sampling $n_{c}+n_{q}$ time steps from a window without replacement. Training tasks are sampled on-the-fly from training windows. For validation and testing, we pre-generate fixed sets of tasks:
  • Validation: 22,178 tasks (batched in 32), with $n_{q}=256$ and $n_{c}$ sampled as in training.
  • Test: 66,586 tasks (batched in 32), with $n_{q}=50$ and $n_{c}$ sampled as in training.